Our engineers where alerted via our network monitoring to a DDoS attack (Distributed Denial of Service) on one of our ESXi servers.
The attack and traffic has been mitigated however for the duration any directly connected services would have seen small amounts of packet loss.
We are monitoring the network and apologise for any inconvenience caused.
We are aware of an issue with our ticket system whereby the main website is not loading. It’s also possible emailed tickets are not being processed.
As we do not host the site internally due to diversity, we are investigating the issue with our 3rd party provider.
If your query is urgent then please contact us via phone.
UPDATE 01 – 12:21
We have been advised the root cause has now been addressed and this was caused by a DNS lookup limitation which has been removed while a long time solution is found.
We are aware of a fault where outbound calls are taking up to 30 seconds to connect. Engineers are currently investigating this and will update you as more information becomes available.
UPDATE01 – 09:45
We are seeing outbound calls starting to connect as normal again. Inbound should also now be working.
UPDATE01 – 10:30
We have requested further details from our wholesaler as to the cause of the issue.
We have just installed our new SSL Certificate for the Plesk control panel at https://easyhttp.co.uk:8443
The certificate expires October 13, 2014 (404 days from today)
Certificate Name matches easyhttp.co.uk
Valid from 29/Aug/2013 to 13/Oct/2014
Issuer AlphaSSL CA – G2
Please note, anyone who gets an error after the 13/09/2013 advising the Certificate is invalid. please flush your browser cache.
We are aware of an issue with outbound SMTP connections being refused and requesting an SSL connection. As we don’t support SSL these are also being refused.
We are taking the service offline at 17:30 to perform a software upgrade to the SMTP service.
*UPDATE 01 @ 17:53**
The upgrade is now complete which we hope will resolve the above issue, However we are also taking this Opportunity to update a few other components.
*UPDATE 01 @ 19:50**
This work is now complete.
We will be upgrading our bonded platform from version 2.19 to 2.32. During this time your router may disconnect a few times, and there may be a period of instability. We expect no more than 30-60 seconds of downtime during this window.
UPDATE 01 @ 23:30
This upgrade is now complete and we are monitoring the service.
We are currently investigating reports of packet loss across the LLU Network. O2 have provided us with the following update:
We are currently investigating a possible problem within our core network, which may cause packet loss and intermittent sessions for your customers. We will let you know as soon as we have any further information.
The above issue may cause issue with the quality of VoIP calls, we will provide more updates as they become available.
Apologies for any inconvenience caused.
UPDATE 2200 07/05/2013
We have received the below update from O2.
The core links are back in service. However, we are still testing the connections and increased latency might be affecting the broadband service in Sheffield and Northampton areas.
We believe the packet loss issue to now be resolved, we will continue to monitor the situation.
Again please accept our apologies for any inconvenience caused today.